Sol Summary – November 13th

MDRS Daily Summary Report for Sol 10

Summary Title: About pancakes.

Mission Status: We had discussed how long time and accuracy is necessary for to do pancakes.:) Sometimes, common circumstances, as prepare a meal and let everything fixed, would consume a great time. Less is more! Some activities had been shared between the crew members. Most of us are available to help each other in regular activities.

Sol Activity Summary: Regular operational activities; preparing meals; each crew member developing their researches; Julio Rezende doing physical activities; Tentative of Carmen and Brandon fix the water heater; EVA of Atila and Camila in front of the MDRS.

Look Ahead Plan: Tomorrow, We are planning an EVA with ATV in the region of Yellow Moon. Julio Rezende will be writing the data related with sustainability.

Anomalies in work:  Nope.

Weather: 7o C, 45o F

Crew Physical Status: Very well.

EVA: The EVA was close to MDRS. Atila and Camila collected some soil samples to a post-analysis.

Reports to be file:

Commanders Report

EVA #6 report

Daily Summary Report

Journalist Report

Biologist Report

Engineering Report

EVA #7 request

GreenHab Report

HSO Report

XSO Report

Support Requested: Nope.

Commander Report – November 13th

6h30m – Awake.

7h50m – 7h50m – 60 minutes walking: 5,6 km or 3,5 miles.

10h20m – 11h – EVA activity walking to collect more soil to a posterior analysis.

11h – Part of our crew is preparing pancakes.

12h – Camila and Brandon tried to fix the water heater, but they didn’t got.

14h30m – The crew is preparing the lunch.

15h – I just spend great part of my time reflecting about what dimensions would be considered about the sustainability study of a Mars analogue habitat. So, the following chart summarizes important key elements to be considered, evaluated in MDRS and propose some improvements about.

Chart 1 – Dimensions of the sustainability to be studied in MDRS

Environmental Sustainability – Activities of environmental education;

– Waste management practices;

– Save water practices;

– Procedures to save energy and use of renewable energy;

– Sustainable Development Goals (SDGs).

Social sustainability – Education related to the research station management;

– Publication of reports;

– Quality of life;

– Equal distribution of activities between genres;

– Occupational Health and Safety Policy;

– Use of social media;

– Code of ethics or code of conduct;

– Social responsibility actions.

Personal sustainability – develop personal resilience strategies to achieve goals;

– maintenance of good mental condition;

– Health;

– team assistance and relationship;

– orientation to activities and tasks;

– Feeling of satisfaction or completude.

Economic sustainability – Operations and management of the simulation;

– Scientific process.

15h30m – Lunch.

16h20m – I walked 1h40m. Today I completed 10 miles, as yesterday. An average of 8 miles by day. This is part of my experiment related with personal sustainability. Apple Watch is providing the info that I supposed had burned more than 1000 calories.

19h – CapComm.

GreenHab Report – November 13th

Sol 10 – 13/11/2017

Green Hab Officer Atila Meszaros

Ambient: heating 50 F

Ambient with window/door open: door opened at 12:32; closed at 17:05

Shade cloth on/off: on

Average temperature:

 

Time 08:56 12:52 17:05 18:44
Temperature inside (°C) 28 38 17 17
Temperature outside (°C) 2 12 10 5
% Humidity 16 16 16 16

Changes to crops:  Crops keep growing! Check the Excel for more details.

Daily water usage for crops: 2 galons

Time(s) of watering for crops: 08:56 and 17:05

Changes to research plants: the amaranthus control keeps growing, mustards too!

Daily watering and amount of water used: 1 ½ L approximately

 

Narrative:

Hi, there, Graeme! Atila in here. I hope you are having an amazing night

The crops from the former experiment in the Green Hab (GH) keep growing. Shannon will have to give them a look because the tomato seedlings are like a pest! I can’t believe how many seedlings are. I hope Shannon give me some instructions regarding this issue. They are growing pretty respectable, though.

I’ve made an excel with all the data of the crops and research plants. It’s only 15 KB, so I’m attaching it to the report.Each square is a pot, I hope it gives you a nice view of the GH.

However, it will be probably easier if I make a Drive like the Excel document with the dates and number of emerging seeds. And this Drive can be used for future crews too. Let me know if you have like a Data Base where I can put this kind of data.

The analog soil is not becoming clay anymore, thanks to a more rigorous water treatment, so I’m waiting for the seedling to show on Wednesday.

I think that’s pretty much all, if you need more information about the GH I will be glad to send it to you.

Thank you for being there!

Have a beautiful night on Earth.

Support/supplies needed: none

Operations Report – November 12th

Crew 182 Operations Report  12Nov2017

SOL: 9

Name of person filing report:  Carmen Atauconcha

Non-nominal systems:

WATER HEATER:  I am going to check the information that you need tonight. I will send the information tomorrow early.

Generator—  It worked the whole night since 5: 00pm of 11Nov2017 until 8:00 am of 12Nov2017.  Turn on at 4:30 pm 12Nov2017.  Operand.

Solar—

READING FROM 12 NOV 2017

TIME SOC READINGS
8:50 am 100%
9:50 am 100%
10:50 am 100%
11:50 am 100%
12:50 am 100%
13:50 pm 98%
14:50 pm 95%
15:26 pm 93%
15:49 pm 89%
16:08 pm 85%
17:40 pm 76%
19:40 pm 75%
20:40 pm 75%

Diesel – 55 %

Propane – 50 psi

Gasoline (5 Gallon containers for ATV) – 0 gallons. In addition, the tank of gasoline inside the ATV’s is almost filled.

Water (trailer) –  0 gallons

Water (static) – 218  gallons

Trailer to Static Pump used –   no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  39.4gallons

Toilet tank emptied:  yes

ATV’s Used: No

Oil Checked:  No

Atv # Fuel Used Gals:  0  gallons

Tires Status: Perfect

# Hours the ATVs were Used today: 0 hours

Deimos used:  No

Hours: 0

Beginning charge:  –

Ending charge:  –

Currently charging: Yes

HabCar used? No

Notes and Comments:
We made bread finally,  it was tasty and we liked it too much.  From another hand, the last night was a little hot. As a result, since tonight we are going to lower the temperature of the heater.
Summary of Hab operations:

We receive the new radios today in the morning.  Also, according to the instructions they were installed in the desk of radios and a couple in the laboratory.  We proved the six radios and they are fine. All of them load and we can hear clearly the messages.   According your question about how many the radios remain are operational, we had 5 remained radios working each one with their own earbud.  In addition, we had three more earbuds; one of them is not working.

According to the toilet, I did what you wrote me almost exactly. The thing is that I did not use the deodorizer pack.  I am sorry for that. I thought that with the digester will be enough.  I am going to empty the toilet tomorrow again, so, I will use the deodorizer pack.

Summary of GreenHab operations:

I will remember that the heater is automatically. Thanks mission support.

Also, inside the greenhab, there is a big blue tub. What is that?

Summary of ScienceDome operations:

We got all the cords off the middle table and put them against the wall.  We worked today until 6:30pm, and there was not light. Therefore, it is necessary the lamps.

Summary of health and safety issues:

While we were trying to cook the lunch my classmate burned his finger. I t was a little little burn. Thanks for the first aid kit, it help us in the situation.

Questions, concerns and requests to Mission Support:

I have a doubt, which is the minimum limit for the static water tank?

We need gasoline for the ATV’S, please.

Sol Summary – November 12th

MDRS Daily Summary Report for sol 9

Summary Title: It’s sunday.

Mission Status: Today looks like a really Sunday. The crew is more relaxed but also focused with their duties. We had a good meal today prepared by Camila. Some of us shared some activities at this time, preparing the lunch. We are a team!

Sol Activity Summary: Keeping a High State of Consciousness in MDRS.

Look Ahead Plan: Tomorrow, We are planning an EVA walking to Lith Cannyon.

Anomalies in work:  Nope.

Weather: 9o C, 49o F

Crew Physical Status: Very well.

EVA: We did an interesting EVA without ATV. We visited an interesting spot walking to the north direction from the MDRS. Distance: 1,2 km. Was collected soil and rock samples. Was possible to do some beautiful images from the area.

Reports to be file: We sent in CapComm.

Support Requested: Nope.

EVA Report – November 12th

EVA #5 Report Sol 9

We began our EVA at 10:40 AM today. We started a bit late because we needed more time to prepare Carmen’s flight suit and test our communication equipment before this EVA. Since we didn’t need to use the rovers today, Julio and Carmen tried the Exploration suits. They didn’t seem particularly more comfortable as compared to the normal Suits. We walked towards the North Ridge area, until I could find a slope on the western slope that was shallow enough for us to climb. Once we found a shallow enough slope, we took samples from different levels of the hill. We took four samples, three from light-dark purple layers of regolith and one from a rusty-colored layer. We took several photographs of the surrounding landscape as well as the sample sites for later reference.

No falls or accidents on this outing. Sample Acquisition was successful! We began recompression at 11:55 PM. Distance covered was about 1.2-1.5 km total. Sample collection and ground covered went much quicker than expected. It’s important to note how quickly the shoulders and upper back can fatigue, even though we were only out for over an hour.

Prepared by Brandon Ferguson

GreenHab Report – November 12th

Sol 9 – 12/11/2017

Green Hab Officer Atila Meszaros

Ambient: heating 50 F

Ambient with window/door open: door opened at 11:32; closed at 17:00

Shade cloth on/off: on

Average temperature:

Time 08:40 13:41 15:36 18:46
Temperature inside (°C) 23 27 21 22
Temperature outside (°C) 2 13 12 8
% Humidity 16 16 16 16

Changes to crops:  Crops keep growing! A lot.

Daily water usage for crops: 2 galons

Time(s) of watering for crops: 08:40 and 15:36

Changes to research plants: first day of germination for amaranthus control! Finally!

Daily watering and amount of water used: 1 ½ L approximately

Narrative:

Hi, Jennifer! Atila in here, yes, like the Hun. I hope you are having an amazing night

The crops from a former experiment that is in the Green Hab (GH) keep growing. There is a gigantic bean seedling! It looks amazing. However, I’m worried about the tomato seedlings, they are too many in the same pot. I hope Shannon give me some instructions regarding this issue.  I elevated the temperature of the GH to 50 F.

Species (common name) Number of bowls Number of seedlings
Tomatoes 25 43
Cucumber 7 3
Melon 8 8
Beans 17 14
Pepper 9 3

Tomorrow as soon as possible I’ll update this numbers. It will be probably easier if I make a Drive with the dates and number of emerging seeds. Let me know if you have like a Data Base where I can put this kind of data.

Respecting my project. Half of the seeds remain at the GH; the other half is in the growing chamber. For all the seeds of the experiment, I’m using almost a Liter.

The project wants to prove that Peruvian native crops can, indeed, germinate and grow on Mars soil. I’m working with quinoa, amaranthus, and mustard (as a control).

The analog soil I’m using is becoming clay, and I’m worry that the seeds are not going to be able to break the soil and emerge. So, I’m going to germinate the seeds on Petri dishes and observe growing behavior on those ones. I already started today,

I think that’s pretty much all if you need more information about the GH I will be glad to send it to you.

Thank you for being there, Jenn!

Have a beautiful night on Earth.

Support/supplies needed: Distilled water (not extremely necessary for now)

Commander Report – Nov 12th

4h – Awake.

4h50m – 7h10m – 140 minutes walking: 12 km or 7,5 miles. Was beautiful to see so many stars in the night. I really like deserts during the night. It is a gift woke up early and have an experience like this.

Today, when I woke up, was 0o or 32oC. Walking in 0 or -2oC is an expression of resilience. When you are focused, there is no pain. This is a characteristic that must be considered in the astronauts experience.

I’m very thankful about the opportunity to prepare myself for this mission and keep the exercises during the simulation in MDRS.

9h – Breakfast of tortillas and cheddar cheese with coffee and

10h10m – 12h – EVA – I considered that We walked from the EVA and returning doing a distance around 2,5 km.  Our crew, always When We are in the decomprehension  chamber, to start an Extravehicular Activities, part of the crew who stays in the Hab plays a music to present an idea of time. The music of decompression was Coming Home of Iron Maiden. Interesting have a music that talks about aerospace.

During the EVA, I used the exploration spacesuit prototype. This spacesuit presents a painful ergonomics. It is hard have movements to collect rocks and soil samples.

13h – I’m doing my commander report. I organized my room and helped in the lunch preparation.

14h – 15h – Lunch.

16h – 17h30m – I took a nap. I was tired about 16 km or 10 miles walking today.

17h30m – From Brazil, I received news about the transmission of the interview of TV Globo related to the research associating Mars, MDRS and Brazilian semiarid. Some images from the visit in MDRS:

http://g1.globo.com/fantastico/edicoes/2017/11/12.html?utm_source=email&utm_medium=share-bar-desktop&utm_campaign=share-bar&#!v/6283988

17h30m – During the afternoon, after the lunch, Atila, Brandon and Camila worked in their research. I took some photos about sent in Crew-photos e-mail.

19h – CapComm.